a family decides to build a small basketball court so thier children can play at home. A concrete pad is to be poured that is 1/2 ft thick, and has dimensions of 15ft by 15 ft. One cubic yard is equivalent to 27ft cubed. How many cubic yards of concrete must be poured?
Answer by mananth(16946)

a family decides to build a small basketball court so thier children can play at home. A concrete pad is to be poured that is 1/2 ft thick, and has dimensions of 15ft by 15 ft. One cubic yard is equivalent to 27ft cubed. How many cubic yards of concrete must be poured?
L= 15ft = 5 yards
W=15ft=5yards
thick = 1/2 ft = 1/2 /(3)= 1/6 yards
...
Volume = V= 5*5*1/6 = 25/6 cu. yards
...
m.ananth@hotmail.ca